Wildhorse Company’s budgeted sales and budgeted cost of goods sold for the coming year are $144,820,000 and $110,160,000, respectively. Short-term interest rates are expected to average 10%. If Wildhorse can increase inventory turnover from its present level of 9 times a year to a level of 12 times per year.

Compute its expected cost savings for the coming year.
Answer #1

As we know

Inventory Turnover = Cost of Goods Sold /Average inventory

110,160,000/Average inventory = 9

Average inventory = 110,160,000/9 = 12,240,000

New inventory Turnover is expected to be 12 times

110,160,000/Average inventory = 12

New Average inventory = 110,160,000/12 = 9,180,000

Cost savings

= (12,240,000 - 9,180,000) * 10%

= 306,000